News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Tuesday 31 August 1999 12.30am EST

The P9 load began executing at 1am on 8/30 and has operated as expected for the past 40 hours. In parallel with P9 execution a series of real-time HRC activities were completed that included first light to HRC with LMC-X1 (another cheer!) and a series of HRC shutter focus measurements. The initial focus was determined using the shutter scans and the values will be available for the P10 load. The initial assessment by the HRC team was that the images look good and the detector is performing well.

A fix to the problem identified with the software commanding of HETG has been approved by the Level 2 board and will be exercised tonight during the radiation passage.

The P9 load will continue to operate until late evening on 9/1.





Return to Status Reports